# Entanglement signatures of a percolating quantum system

### v1

@@ Abstract
Outlook: Quantum entanglement and its measures have taken a defining role in deciphering the nature of quantum phases. In particular, the nature of low energy excitations are often equated with whether the bipartite entanglement entropy follows area law, or has logarithmic corrections. In this work, we revisit various measures of quantum entanglement in the context of percolation disorder in free fermionic lattice Hamiltonians. We find that percolation disorder inherently generates extensive degeneracies, which gives rise to subtleties in standard bipartite entanglement entropy. It is then important to either break the massive degeneracies by putting infinitesimal disorder which leads to the $S_{\text{quen}}$, or otherwise investigate quantities such as $S_{\text{typ}}$ and $S_{\text{ann}}$ which includes physics of the degenerate manifold. We uncover that such measures have contributions from both fermionic correlations and geometrical aspects. While $S_{\text{quen}}$ generically follow area law, $S_{\text{typ}}$ and $S_{\text{ann}}$ are volume law in character. These quantities can, in turn, be estimated from properties of the clusters, which either cut the entanglement bi-partition or contribute to the zero-mode degeneracies. Interestingly, the entanglement measure of the largest cluster can capture even the classical percolation threshold in two dimensions.
@@
While in this work, we do not propose any experimental setups to measure such entanglement signatures, finding realistic proposals [79, 80, 81, 16] in this direction would be interesting to pursue. Similarly, we have focused in this study exclusively on bipartite entanglement entropy, given its relevance to quantum condensed matter systems. In quantum information, various other measures have been pursued to investigate phases and phase transitions [82, 83]. A comprehensive investigation of these, in regard to percolation disorder, is another prospective study. Finally, investigating this physics in both symmetry-protected topological systems and topologically ordered systems will be an exciting future direction.
@@ I Entanglement between two free spins
To understand the notion of typical and annealed entanglement in contrast with standard bipartite entanglement entropy, let us consider two free spins $A$ and $B$ as shown in Fig. S5(a). A general choice of the valid many-body wavefunction of the system will be,
@@
$\hat{\rho}_{A}$ $=$ $\text{Tr}_{B}(\hat{\rho})=\text{Tr}_{B}\big{(}\langle\psi|\psi\rangle\big{)}$ (S3)
@@
Now using $\hat{\rho}_{A}$ one can numerically calculate bipartite entanglement entropy $S_{A}(\hat{\rho}_{A})=\text{Tr}\big{(}{\hat{\rho}_{A}\log\hat{\rho}_{A}}\big{)}$ and average it over numerous Gaussian ensembles of wavefunction coefficients {$a,b,c,d$} to find typical entanglement measure $S_{\text{typ}}=\langle S_{A}(\hat{\rho}_{A})\rangle\simeq 0.33$. We also define annealed entanglement as $S_{\text{ann}}=S_{A}(\langle\hat{\rho}_{A}\rangle)=\text{Tr}\big{(}{\langle\hat{\rho}_{A}\rangle\log\langle\hat{\rho}_{A}\rangle}\big{)}$. To calculate this, one first takes ensemble averaged (this average is also taken over wavefunction coefficients as described earlier) partial density matrix $\langle\hat{\rho}_{A}\rangle$ which turns out to be $\frac{1}{2}I_{2}$ here, and then find entanglement entropy to get $S_{\text{ann}}=\log 2$.
Figure S5: (a) Two free spin $A$ and $B$. (b) A four-site toy model with only two sites ($1$ and $4$) are connected via fermionic hopping of strength $t$. The bipartition used for entanglement calculation is shown in red-dashed line. (c) Different contributions in $S_{\text{typ}}$ for the four-site toy example. Contribution from fermionic correlation $S_{\text{corr}}$ is always $\log 2$. For complex random coefficients, the geometric contribution $S_{\text{geom}}$ converges to $S_{\text{geom}}(\mathbb{C})=\frac{1}{2}$ with increasing $N_{R}$ (number of random pure state ensembles), and for real choices it converges to $S_{\text{geom}}(\mathbb{R})=(2\log 2-1)$.
@@ II Four site fermionic toy example
In order to illustrate how typicality in entanglement measure appears in a fermionic model, consider the system where two spinless fermions hops (with hopping strength $t=-1$) on a four-site structure as shown in Fig. S5(b). While the single-particle spectrum is straight-forwardly given by {±$\{\pm 1,0,0\}$, the many-body ground state is not unique but two-fold degenerate. This can be simply seen since, to minimize the energy, while one fermion resides in a delocalized state between sites 1 and 4, another fermion can either be on either site 2 or 3. The two valid ground-state wavefunctions of the system of the system are
@@
where $|\Omega\rangle$ is the fermionic vacuum and the entanglement entropy across the partition is clearly $\log 2$ given the delocalized fermion. However, any general wavefunction of the form
@@
is also a valid wavefunction of the system. Given the coefficients are randomly drawn from a Gaussian distribution and follow wavefunction normalisation constraint, they form a random pure state similar to what we discussed in I. This leads to similar considerations of typical entanglement $S_{\text{typ}}$ as described in the main text. However, now it has components arising from the fermionic correlations within the connected cluster across the partition (the delocalized state between site 1 and 4) and contributions due to a random choice of coefficients $\phi_{1}$ and $\phi_{2}$. In general,
@@
where $S_{\text{corr}}$ captures the fermionic correlations in connected regions and $S_{\text{geom}}$ arises out of typicality due to degeneracy, often arising out of disconnected sites. Therefore, the latter has information about the geometrical properties of the network. In the present example, while $S_{\text{corr}}=\log 2$, $S_{\text{geom}}$ depends on the choices of the wavefunction coefficients ($\phi_{1}$ and $\phi_{2}$) used in the formation of random pure states, leading to the symmetry classification of the typical entanglement [55]. In class A [58], the coefficients are complex, and the analytical result shows $S_{\text{geom}}(\mathbb{C})=\frac{1}{2}$, while in class AI [58] where the coefficients are real, it is evaluated to be $S_{\text{geom}}(\mathbb{R})=(2\log 2-1)$. In Fig. S5(c), we show that our numerical results converge towards analytical results with the increasing number of pure ensembles ($N_{R}$) used in the calculation of $S_{\text{typ}}$.
@@ III Zero-dimensional fermionic problem
An all-to-all connected fermionic hopping network hosts $N_{0}=N-1$ number of zero energy states for a $N$ site system. So, a typical many-body wavefunction of the system will be a random pure state made out of these degenerate zero energy states. Here, we choose the wavefunction coefficients to be real, restricting it to class AI [55]. Now, $S_{\text{typ}}$ at half filling $f=\frac{1}{2}$, and for equal bipartition, typical entanglement contain a small fermionic correlation part and a huge geometric contribution because of the massive degeneracy in the spectrum. While the geometric part follows the symmetry classification result [55], the fermionic part is captured in $S_{\text{quen}}$, evaluated using the degeneracy breaking protocol as discussed in the main text. Now for any general filling $f$ ($10<f<$ 1) with equal bipartition, the available phase space volume is $f(1-f)$; consequently, the effective degenerate Hilbert space dimension becomes $\sim 2^{af(1-f)N_{0}}$, where $a$ is a dimensionless constant. Then, using Page’s result [48], we empirically find the typical entanglement in the leading order to be,
$S_{\text{typ}}(f)=S_{\text{quen}}(f)+4f\big{(}1-f\big{)}\big{(}N_{0}-1\big{)}\big{(}\log 2-\frac{1}{2}\big{)},$ (S9)
so that it remains consistent with the result at half filling ($S_{\text{typ}}=S_{\text{quen}}+\big{(}N_{0}-1\big{)}\big{(}\log 2-\frac{1}{2}\big{)}$). In Fig. S6, we show the numerical values as a function of filling ($f$), which match well with our empirical formula.
IV Lattice animals, their entanglement and number of zeromodes
In the case of square lattice percolation, the quenched entanglement $S_{\text{quen}}$ can be captured using the lattice animals and their contributions to entanglement given a bipartition. For a lattice animal of bond size $k$ (i.e. number of bonds in that lattice animal is $k$) and perimeter $t$, we calculate the bipartite entanglement entropy $S_{\text{animal}}^{b}$ with the partition dissecting the animal in a certain way. Now if $P_{\text{animal}}^{b}$ is the probability that the different animals of the same {$k,t$} appear on that given partition and give rise to the same $S_{\text{animal}}^{b}$, then for a square lattice of linear size $L$,
$N_{0}=\sum_{k,t}n_{0}(k,t)g(k,t)p^{k}(1-p)^{t}L^{2},$ (S11)

### v2

@@ Abstract
Outlook: Quantum entanglement and its measures have taken a defining role in deciphering the nature of quantum phases. In particular, the nature of low-energy excitations are often equated with whether the bipartite EE follows area law or has logarithmic corrections. In this work, we revisit various measures of quantum entanglement in the context of percolation disorder in free fermionic lattice Hamiltonians. We find that percolation disorder inherently generates extensive degeneracies, which gives rise to subtleties in standard bipartite EE. It is then important to either break the massive degeneracies by putting infinitesimal disorder which leads to the $S_{\text{quen}}$, or otherwise investigate quantities such as $S_{\text{typ}}$ and $S_{\text{ann}}$ which includes physics of the degenerate manifold. We uncover that such measures have contributions from both fermionic correlations and geometrical aspects. While $S_{\text{quen}}$ generically follow area law, $S_{\text{typ}}$ and $S_{\text{ann}}$ are volume law in character. These quantities can, in turn, be estimated from the properties of the clusters, which either cut the entanglement bipartition or contribute to the zero-mode degeneracies. Interestingly, the entanglement measure of the largest cluster can capture even the classical percolation threshold in two dimensions. While we have restricted our investigation to three quantities $S_{\text{typ}},S_{\text{ann}}$ and $S_{\text{quen}}$, it would be worthwhile to quantify the amount of classical and quantum correlations in these systems. In physical systems where a perturbative quenched disorder is inherent, it is expected that $S_{\text{quen}}$ represents a more meaningfully observable quantity than $S_{\text{typ}}$. Similarly, given in estimation of $S_{\text{ann}}$ complete density matrix is averaged, the annealed EE contains both classical and quantum correlations. In this context, quantities such as mutual information and entanglement negativity may be of interest.
@@
While in this work, we do not propose any experimental setups to measure such entanglement signatures, finding realistic proposals [97, 98, 99, 16] in this direction would be interesting to pursue. Similarly, we have focused in this study exclusively on bipartite EE, given its relevance to quantum condensed matter systems. Various other measures have been pursued in quantum information to investigate phases and phase transitions [100, 101]. A comprehensive investigation of these, in regard to percolation disorder, is another prospective study. Finally, investigating this physics in both symmetry-protected topological systems and topologically ordered systems will be an exciting future direction.
@@ I Entanglement between two free spins
To understand the notion of typical and annealed entanglement in contrast with standard bipartite entanglement entropy, let us consider two free spins $A$ and, $B$ as shown in Fig. S5(a). A general choice of the valid many-body wavefunction of the system will be,
@@
$\hat{\rho}_{A}$ $=$ $\text{Tr}_{B}(\hat{\rho})=\text{Tr}_{B}\big{(}|\psi\rangle{\langle\psi|}\big{)}$ (S3)
@@
Now using, $\hat{\rho}_{A}$ one can numerically calculate bipartite entanglement entropy $S_{A}(\hat{\rho}_{A})=\text{Tr}\big{(}{\hat{\rho}_{A}\ln\hat{\rho}_{A}}\big{)}$ and average it over numerous Gaussian ensembles of wavefunction coefficients {$a,b,c,d$} to find typical entanglement measure $S_{\text{typ}}=\langle S_{A}(\hat{\rho}_{A})\rangle\simeq 0.33$. We also define annealed entanglement as $S_{\text{ann}}=S_{A}(\langle\hat{\rho}_{A}\rangle)=\text{Tr}\big{(}{\langle\hat{\rho}_{A}\rangle\ln\langle\hat{\rho}_{A}\rangle}\big{)}$. To calculate this, one first takes ensemble averaged (this average is also taken over wavefunction coefficients as described earlier) partial density matrix $\langle\hat{\rho}_{A}\rangle$ which turns out to be $\frac{1}{2}I_{2}$ here, and then find entanglement entropy to get $S_{\text{ann}}=\ln 2$.
Figure S5: (a) Two free spin $A$ and $B$. (b) A four-site toy model with only two sites ($1$ and $4$) are connected via fermionic hopping of strength $t$. The bipartition used for entanglement calculation is shown in red-dashed line. (c) Different contributions in $S_{\text{typ}}$ for the four-site toy example. Contribution from fermionic correlation $S_{\text{corr}}$ is always $\ln 2$. For complex random coefficients, the geometric contribution $S_{\text{geom}}$ converges to $S_{\text{geom}}(\mathbb{C})=\frac{1}{2}$ with increasing $N_{R}$ (number of random pure state ensembles), and for real choices it converges to $S_{\text{geom}}(\mathbb{R})=(2\ln 2-1)$.
@@ II Four site fermionic toy example
To illustrate how typicality in entanglement measure appears in a fermionic model, consider the system where two spinless fermions hops (with hopping strength $t=-1$) on a four-site structure as shown in Fig. S5(b). While the single-particle spectrum is straight-forwardly given by {±$\{\pm 1,0,0\}$, the many-body ground state is not unique but two-fold degenerate. This can be seen because, to minimize the energy, while one fermion resides in a delocalized state between sites 1 and 4, another fermion can be on either site 2 or 3. The two valid ground-state wavefunctions of the system are
@@
where $|\Omega\rangle$ is the fermionic vacuum and the entanglement entropy across the partition is clearly $\ln 2$ given the delocalized fermion. However, any general wavefunction of the form
@@
is also a valid wave function of the system. Given the coefficients are randomly drawn from a Gaussian distribution and follow wavefunction normalization constraint, they form a random pure state similar to what we discussed in I. This leads to similar considerations of typical entanglement $S_{\text{typ}}$ as described in the main text. However, now it has components arising from the fermionic correlations within the connected cluster across the partition (the delocalized state between site 1 and 4) and contributions due to a random choice of coefficients $\phi_{1}$ and $\phi_{2}$. In general,
@@
where $S_{\text{corr}}$ captures the fermionic correlations in connected regions and $S_{\text{geom}}$ arises out of typicality due to degeneracy, often arising out of disconnected sites. Therefore, the latter has information about the geometrical properties of the network. In the present example, while $S_{\text{corr}}=\ln 2$, $S_{\text{geom}}$ depends on the choices of the wavefunction coefficients ($\phi_{1}$ and $\phi_{2}$) used in the formation of random pure states, leading to the symmetry classification of the typical entanglement [69]. In class A [75], the coefficients are complex, and the analytical result shows $S_{\text{geom}}(\mathbb{C})=\frac{1}{2}$, while in class AI [75] where the coefficients are real, it is evaluated to be $S_{\text{geom}}(\mathbb{R})=(2\ln 2-1)$. In Fig. S5(c), we show that our numerical results converge towards analytical results with the increasing number of pure ensembles ($N_{R}$) used in the calculation of $S_{\text{typ}}$.
@@ III Zero-dimensional fermionic problem
An all-to-all connected fermionic hopping network hosts $N_{0}=N-1$ number of zero energy states for a $N$ site system. So, a typical many-body wavefunction of the system will be a random pure state made out of these degenerate zero energy states. Here, we choose the wavefunction coefficients to be real, restricting it to class AI [69]. Now, $S_{\text{typ}}$ at half filling $f=\frac{1}{2}$, and for equal bipartition, typical entanglement contains a small fermionic correlation part and a substantial geometric contribution because of the massive degeneracy in the spectrum. While the geometric part follows the symmetry classification result [69], the fermionic part is captured in $S_{\text{quen}}$, evaluated using the degeneracy breaking protocol as discussed in the main text. Now, for any general filling $f$ ($10<f<$ 1) with equal bipartition, the available phase space volume is $f(1-f)$; consequently, the effective degenerate Hilbert space dimension becomes $\sim 2^{af(1-f)N_{0}}$, where $a$ is a dimensionless constant. Then, using Page’s result [62], we empirically find the typical entanglement in the leading order to be,
$S_{\text{typ}}(f)=S_{\text{quen}}(f)+4f\big{(}1-f\big{)}\big{(}N_{0}-1\big{)}\big{(}\ln 2-\frac{1}{2}\big{)},$ (S9)
so that it remains consistent with the result at half filling ($S_{\text{typ}}=S_{\text{quen}}+\big{(}N_{0}-1\big{)}\big{(}\ln 2-\frac{1}{2}\big{)}$). In Fig. S6, we show the numerical values as a function of filling ($f$), which match well with our empirical formula.
IV Mutual information in one-dimensional percolation
In general, mutual information (MI) is a good measure of entanglement between two subsystems $A$ and $B$ when the full system is in mixed state [71]. Since finding annealed entanglement entropy ($S_{\text{ann}}$) in our study is associated with the mixedness of the full system, here we intend to show the behavior of MI in the percolated tight-binding chain. With the same bipartition protocol used in the main text, MI between $A$ and $B$ is,
$I_{A:B}=S^{A}_{\text{ann}}+S^{B}_{\text{ann}}-S^{AB}_{\text{ann}},$ (S10)
where, the annealed entanglement entropy of the full system $S^{AB}_{\text{ann}}\neq 0$ due to mixed character of the correlator matrix for the full system at half filling. In Fig. S7, the behavior of disorder averaged MI ($\tilde{I}_{A:B}$) with bond occupation probability $p$ is shown for our one-dimensional percolation model. In $p\rightarrow 0$ limit, $S^{A}_{\text{ann}}+S^{B}_{\text{ann}}\approx S^{AB}_{\text{ann}}$. While the behavior shows reduction of large volume law (due to zeromodes) contribution in $S_{\text{ann}}$ for small $p$ values, a sharp rise appears near the classical percolation threshold $p=p_{c}=1$, where the lattice gets connected. This is similar to $\tilde{S}_{\text{quen}}$ (see Fig. S7), which is expected because, after the removal of zeromode contribution, only the electronic correlation part reflects in MI.
Figure S7: Disorder averaged MI $\tilde{I}_{A:B}$ with percolation probability $p$ for the free fermionic chain of length $L=200$. The bipartition protocol is the same as the main text. The disorder average is taken over $500$ configurations, and for each configuration, $10^{3}$ random pure states are chosen for averaging the correlator matrix. Note the similarity in behaviour with $\tilde{S}_{\text{quen}}$ (same data as shown in the main text for $L=200$).
V Lattice animals, their entanglement and number of zeromodes
@@ 00 $q^{4}$ $0,4$ $1$ $1$
In the case of square lattice percolation, the quenched entanglement $S_{\text{quen}}$ can be captured using the lattice animals and their contributions to entanglement given a bipartition. For a lattice animal of bond size $k$ (i.e., the number of bonds) and perimeter $t$, we calculate the bipartite entanglement entropy $S_{\text{animal}}^{b}$ with the partition dissecting the animal in a certain way. Now if $P_{\text{animal}}^{b}$ is the probability that the different animals of the same {$k,t$} appear on that given partition and give rise to the same $S_{\text{animal}}^{b}$, then for a square lattice of linear size $L$,
$N_{0}=\sum_{k,t}n_{0}(k,t)g(k,t)p^{k}(1-p)^{t}L^{2},$ (S12)